If you have 15.0 g Na, how many grams of O2 are required for reaction?
1
Expert's answer
2020-10-02T14:23:26-0400

STEP I : BALANCED EQUATION AND MOLES OF Na

2Na(s) + O2(g) "\to" 2Na2O(s)


Moles of Na in 15.0g "= \dfrac{mass}{RAM} = \dfrac{15.0g}{23g/mol}"


"=0.652moles"


STEP II : MOLES AND MASS OF O2


From the equation, mole ratio of Na:O2 = 1:1


Moles of oxygen "=\dfrac{1}{2}x0.652mol = 0.326moles"


Mass of oxygen = Moles x Molar Mass

= 0.326mol x 16g/mol

= 5.22g


Mass of oxygen